6. Solve the given initial value problem, with y0 = 2 and y0) = 5: y - 6y + 5y = 20t +1
Answer #1

We have y - by + 5y = 20t + 1 The comes bonding auxiliary equation e 8²-68+5=0 & 2-58-&75 = 0 rls-5) -61-5)=0 (2-1262-5) =Hence the general solution is y(t) = y(t) + Yp (t) cettCq e5t +46+5 » f(t) = Get + 56ge5t + 4 Given. y 601 = 2 C + C + = 2 a